A 17 2-g biece of Styrafoam carries a net charge of 0 658 C and is suspended in equlibriam above the coester of a large, hortonta shee of plasic tht has a uniform charge density on its surface What unit ares on the plastic sheet?
Answer #1

When the styrofoam is suspended, the gravitational force is balanced by the Electric field produced by the sheet below, Hence we can write,

F_g = F_E Rightarrow mg = qE

it9

(17.2 × 10-3 ) (9:81) = 2.56 × 105 N/ 0.658 × 10-6

The electric field produced by the sheet is given by,

E= 2E0

Substituting values we get,

>-= 2.56 × 105 N/C 2Eo

Rightarrow sigma = 4.53 imes 10^{-6} , C/m^2 = 4.53 , mu C/m^2